what is the linear equation of -3x+2=0
0=5x+3
0=-2x+4
-x+1+0
i really need help with these

Answer by edjones(8007)

in order to solve for x you must have x on one side of the equal sign and everything else on the other. in order to do that you add, subtract, multiply or divide any positive or negative number by each side until you get the job done.
-3x+2=0
-3x=-2 subtract 2 from each side.
x=2/3 divide each side by -3.
.
Ed